# Transport, Variational Inference and Diffusions: with Applications to Annealed Flows and Schrödinger Bridges

### v1

@@ 4.1 Iterative proportional fitting (IPF) and the EM algorithm
$\phi_{n+1}$ $=\operatorname*{arg\,min}_{\phi}D_{\mathrm{KL}}(\overrightarrow{{\mathbb{P}}}^{\mu,f+\sigma^{2}\nabla\phi},\overleftarrow{{\mathbb{P}}}^{\nu,f+\sigma^{2}\nabla\theta_{n+1}})$
@@
The key observation in the proof of Proposition 4.1 is that replacing forward-$D_{\mathrm{KL}}$ by reverse-$D_{\mathrm{KL}}$ in one or both of (24a) and (24b) does not – in theory – change the sequence of minimisers. In practice, however, optimisation with respect to forward-$D_{\mathrm{KL}}$ and backward-$D_{\mathrm{KL}}$ encourages moment-matching and mode-seeking behaviour, respectively, and so an alternating scheme as defined in (4.1) might present a suitable compromise. For comparisons of IPF using different directions of $D_{\mathrm{KL}}$ we refer the reader to Vargas et al. (2021a); Zhang et al. (2014).
4.2 Optimality conditions and HJB-regularisers
As per Section 4.1, IPF resolves the nonquniqueness in minimising $\mathcal{L}_{D}(\phi,\theta)$ by performing the coordinate-wise updates (4.1) starting from an initialisation informed by the Schrödinger prior. On the basis of this observation, the joint updates $(\phi_{n+1},\theta_{n+1})\leftarrow(\phi_{n},\theta_{n})-h\nabla_{\phi,\theta}\mathcal{L}_{D}(\phi,\theta)$ suggest themselves, in the spirit of VAEs (Kingma et al., 2019) and as already proposed by Neal and Hinton (1998). However, as is clear from the introduction, the limit $\lim_{n\rightarrow\infty}(\phi_{n},\theta_{n})$, can merely be expected to respect the marginals in (6), and no optimality in the sense of (23) is obtained. As a remedy, we recall the following result:
Proposition 4.3 (Mean-field game formulation).

### v6

@@ E.4 Drift based EM
$\phi_{n+1}$ $=\operatorname*{arg\,min}_{\phi}D_{\mathrm{KL}}(\overrightarrow{{\mathbb{P}}}^{\mu,f+\sigma^{2}\nabla\phi},\overleftarrow{{\mathbb{P}}}^{\nu,f+\sigma^{2}\nabla\theta_{n+1}})$ (73)
@@
Proof.
We can show via the $D_{\mathrm{KL}}$ chain rule and the disintegration theorem (Léonard, 2014b) that the above is minimised when $\theta$ satisfies $\overleftarrow{{\mathbb{P}}}^{\nu,f+\sigma^{2}\nabla\theta}=\overrightarrow{{\mathbb{P}}}^{\mu,f+\sigma^{2}\nabla\phi_{n-1}}\frac{{\mathrm{d}}\nu}{{\mathrm{d}}\rho_{T}^{\mu,f+\sigma^{2}\nabla\phi_{n-1}}}$ which corresponds to $\nabla\theta_{n}=\sigma^{2}\nabla\phi_{n-1}-\sigma^{2}\nabla\ln\rho_{t}^{\mu,f+\sigma^{2}\nabla\phi_{n-1}}$ following Observation 1 in Vargas et al. (2021a). Similarly as per Proposition 3.1 the results will follow for the even half bridges.
EM initialisation:
E.5 HJB-Regularizers
As per Section 3.1, IPF resolves the nonquniqueness in minimising $\mathcal{L}_{D}(\phi,\theta)$ by performing the coordinate-wise updates (19) starting from an initialisation informed by the Schrödinger prior. On the basis of this observation, the joint updates $(\phi_{n+1},\theta_{n+1})\leftarrow(\phi_{n},\theta_{n})-h\nabla_{\phi,\theta}\mathcal{L}_{D}(\phi,\theta)$ suggest themselves, in the spirit of VAEs (Kingma et al., 2019) and as already proposed in this setting by Neal & Hinton (1998). However, as is clear from the introduction, the limit $\lim_{n\rightarrow\infty}(\phi_{n},\theta_{n})$, can merely be expected to respect the marginals in (6), and no optimality in the sense of (17) is expected. As a remedy, we present the following result:
Proposition E.3.
For $\lambda>0$, a divergence $D$ on path space, and $\phi,\psi\in C^{1,2}([0,T]\times\mathbb{R}^{d};\mathbb{R})$, let
$\mathcal{L}$ (${}^{\mathrm{Schr}}(\phi,\theta):=D(\overrightarrow{{\mathbb{P}}}^{\mu,f+\sigma^{2}\nabla\!\phi},\!\!\overleftarrow{{\mathbb{P}}}^{\nu,f+\sigma^{2}\nabla\theta})+\lambda\mathrm{Reg}(\phi),$ (74)
where $\mathrm{Reg}(\phi)=0$ if and only if the HJB-equation $\partial_{t}\phi+f\cdot\nabla\phi+\frac{\sigma^{2}}{2}\Delta\phi+\tfrac{\sigma^{2}}{2}|\nabla\phi|^{2}=0$ holds. Then $\mathcal{L}_{\mathrm{Schr}}(\phi,\theta)=0$ implies that the drift $a_{t}:=\sigma^{2}\nabla\phi_{t}$ solves (17).
The proof rests on an optimal control reformulation of the Schrödinger problem (see Appendix E), identifying the HJB-equation as the missing link that renders joint minimisation of (74) theoretically sound for solving (17). The loss in (74) has two important benefits compared to standard IPF. First, it circumvents the need for the sequential updates used in IPF, thereby simplifying and speeding up the optimisation procedure. Second, it enforces the Schrödinger prior drift $f$ directly, rather than recursively via eq. (18a), (18b). This prevents the prior from being forgotten, as is usually the case in regular IPF. In Appendix E.5, we detail possible constructions of $\mathrm{Reg}(\phi)$, discuss relationships to previous work, and evaluate the performance of the suggested approach in numerical experiments.
This result can be found in Chen et al. (2021, Proposition 5.1), for instance, but since it is relevant to the connections pointed out in Remark 13 below, we present an independent proof:
Proposition E.4 (Mean-field game formulation).